# Extreme magnetic fields around black holes

### v1

@@
Abstract
Recent results of the event horizon-scale images of M87* and Sagittarius A* from the Event Horizon Telescope Collaboration show that strong magnetic fields are likely present around the central black holes (BHs) in these sources. Magnetically arrested disks (MADs), the end stage of magnetic flux saturation around BHs, are especially rich in horizon-scale physics due to the presence of powerful jets and magnetic flux eruptions that provide significant feedback on the accretion mechanism. Here, we present an overview of our current knowledge about the magnetic field evolution in numerical simulations of accreting BHs, focusing on jet launching, black hole-interstellar medium feedback, and black hole imaging of MADs. We find that misaligned MAD accretion flows seemingly exhibit jet ejection cycles that could produce flaring states in radio-quiet active galactic nuclei. Further, we show that advances in horizon-scale interferometric telescopes could identify disk misalignment by imaging the disk-jet connection region.
@@ 1 Introduction
While the idea of magnetically dominated accretion flows has been around for at least 50 years BK1974 , this mode of accretion, dubbed magnetically arrested disks (MADs), truly became a subject of interest over the past 2 decades or so Igumenshchev2003 ; Narayan2003 . Part of the interest stemmed from advances in computational development of 3D general relativistic magnetohydrodynamic (GRMHD) codes that enabled self-consistent simulations of black hole (BH) accretion. On the observational front, the multitude of Sagittarius A* (or Sgr A*) X-ray flares captured by the Chandra observatory Baganoff2001 ; Nowak2012 ; Haggard2019 provided indications of particle acceleration, possibly driven by reconnection of magnetic fieldlines close to the BH Markoff2001 ; Yuan2003 . Figure 1 shows that our standard picture of magnetorotational instability (MRI)-driven accretion produces X-ray and near-infrared (nIR) flares too weak to explain some of Sgr A*’s brightest flares Do2019 ; Haggard2019 , suggesting the need for stronger magnetic fields. More recently, the polarised horizon-scale EHT images of M87 and Sgr A* also favour strong vertical magnetic fields EHTpaperVIIIM872021 ; EHTpaperVIIISgrA2024 . Indeed, the MAD state could be a natural outcome of starved BHs (i.e., accreting at highly sub-Eddington rates), which is the case for both Sgr A* and M87. There is seemingly an abundant supply of poloidal magnetic field either by advecting $\mu$Gauss-level magnetic fields from the interstellar medium (ISM) Narayan2003 ; Haverkorn2013 or from magnetised stellar winds (as thought to occur in Sgr A*; Ressler2020 ), or by in-situ generation of poloidal flux via dynamo processes in the disk Liska2020 , favouring the emergence of the MAD state. Motivated by these advances in theory, numerics and observations, we discuss the many features of MAD accretion flows and their outflows, with a focus on recent results of misaligned MAD flows from Chatterjee et al. Chatterjee2023tiltedMADs .
@@ 2 Revisiting magnetically arrested accretion flows
[…]

### v2

@@
11institutetext: Koushik Chatterjee 22institutetext: Institute for Research in Electronics and Applied Physics, University of Maryland, 8279 Paint Branch Drive, College Park, MD 20742, USA, 22email: kchatt@umd.edu 33institutetext: Matthew Liska 44institutetext: Center for Relativistic Astrophysics, Georgia Institute of Technology, Howey Physics Bldg, 837 State St NW, Atlanta, GA 30332, USA, 44email: mliska3@gatech.edu 55institutetext: Alexander Tchekhovskoy 66institutetext: Center for Interdisciplinary Exploration & Research in Astrophysics (CIERA), Physics & Astronomy, Northwestern University, Evanston, IL 60202, USA, 66email: atchekho@northwestern.edu 77institutetext: Sera Markoff 88institutetext: Anton Pannekoek Institute for Astronomy, University of Amsterdam, Science Park 904, 1098 XH Amsterdam, The Netherlands, 88email: S.B.Markoff@uva.nl 99institutetext: Ramesh Narayan 1010institutetext: Black Hole Initiative at Harvard University, 20 Garden Street, Cambridge, MA 02138, USA, 1010email: rnarayan@cfa.harvard.edu
Extreme magnetic fields around black holes
[…]
   and
[…]
@@ 1 Introduction
[…]
@@ 2 Revisiting magnetically arrested accretion flows
We focus on three quantities: the accretion rate $\dot{M}=-\iint\rho u^{r}dA_{\theta\varphi}$, the poloidal magnetic flux $\Phi_{\rm BH}=(1/2)\iint|B^{r}|dA_{\theta\varphi}$ calculated at the event horizon, and the total outflow power $P=\dot{M}c^{2}-\dot{E}$. For these quantities, we require the gas density $\rho$, the radial component of the gas 4-velocity $u^{r}$, the radial component of the 3-magnetic field vector $B^{r}$ and the total energy flux in the radial direction $\dot{E}=\iint T^{r}_{t}dA_{\theta\varphi}$, defined using the stress-energy tensor $T^{\mu}_{\nu}$ and the area element $dA_{\theta\varphi}=\sqrt{-det(g_{\mu\nu})}d\theta d\varphi$. We assume natural units $G=c=1$, which implies that our characteristic length-scale, the gravitational radius $r_{\rm g}=GM/c^{2}$, and the corresponding light-crossing time-scale $t_{\rm g}=r_{\rm g}/c$, are normalised to the BH mass $M$. For our simulations, we can set $M=1$. Further, we can normalise the magnetic flux and the outflow power with the time-averaged accretion rate $\langle\dot{M}\rangle$, giving us the dimensionless magnetic flux $\phi_{\rm BH}=\Phi_{\rm BH}/\sqrt{\langle\dot{M}\rangle}$ and the outflow efficiency $\eta=P/\langle\dot{M}\rangle$ (note that $r_{\rm g}$ and the speed of light $c$ are unity according to our units). We refer the reader to Porth et al. Porth2019 for more details about these quantities.